OpenStudy (anonymous):

A 1250 kg hippo slides down a mud covered hill inclined at an angle of 18 degrees to the horizontal. If the coefficient of sliding friction between the hippo and the mud is 0.0900, what force of friction impedes the hippo's motion down the hill?

OpenStudy (turingtest):

Here's the free-body diagram:and we have\[F_g=mg\]\[F_f=\mu N\]now we need a little trigonometry to get N...

OpenStudy (turingtest):

We can see that the force into the surface is\[F_g\sin\theta=N\]which will have the same magnitude and opposite direction as the normal force, N. Since for this problem we need only concern ourselves with the magnitude of these vectors we can write the force of friction as\[F_f=\mu N=\mu F_g\sin\theta=\mu mg\sin\theta\]Plug in and solve.